Please help me solve this equation by using substitution. Here is the problem...
1.2x-y=-3
x+y=9
Below is the work that I did to attempt the problem, but i know its not correct.
x+y=9
x=9
(9)2x=18
x=18
18-y=-3
-18
y=-21
18-21=-3

Answer by stanbon(75887)

1.2x-y=-3
x+y=9
------
Add the left sides; Add the right sides.
--------------
2.2x = 6
Solve for "x":
x = 2.73
-------
Solve for "y":
x + y = 9
2.73 + y = 9
y = 6.27
---------
Answer: (2.73,6.27)
======================
Cheers,
Stan H.
==================